# A positive meshless finite difference scheme for scalar conservation laws with adaptive artificial viscosity driven by fault detection

### v1

@@ 4 Adaptive fault-based viscosity
For each ${\bf z}\in X$, let $X_{\bf z}:=\{{\bf x}_{i},\,i=1,...,n_{\bf z}\}\subseteq X$ be the set of $n_{\bf z}$ points closest to ${\bf z}$. We define the fault indicator
$I_{\Delta}({\bf z},X_{\bf z}):=\frac{|\hat{\Delta}f({\bf z})|}{\sum_{i=1}^{n_{\bf z}}|w_{i}|\,\|{\bf x}_{i}-{\bf z}\|^{2}_{2}}\,,$ (28)
@@
$\hat{\Delta}f({\bf z}):=\sum_{i=1}^{n_{\bf z}}w_{i}f({\bf x}_{i})$ (29)
is the $\ell_{2}$-minimal differentiation formula approximating the Laplacian $\Delta f({\bf z})$ with polynomial exactness order $q=3$ and minimizing the seminorm $|{\bf w}|_{2,3}$.
@@
$I_{\Delta}({\bf z},X_{\bf z})\leq|f|_{1,1,\Omega_{\bf z}},$ (30)
with $\Omega_{\bf z}$ as in (22). Although (30) is stated in Bracco et al. [2019] only for the bivariate case, it is easy to extend it to any number of variables $d$. Indeed, this follows from the fact that Theorem 3 in Bracco et al. [2019] holds for any $d$ with the same proof. It is remarkable that the estimate (30) does not contain any unknown constants and therefore allows for comparison between different ${\bf z}$ independently of how irregular the local node set $X_{\bf z}$ is. Therefore, a large value of $I_{\Delta}({\bf z},X_{\bf z})$ indicates that $f$ is likely to be discontinuous in $\Omega_{\bf z}$, or at least its gradient to change abruptly there, the behavior expected from the solution of a conservation law in the vicinity of a shock front. Moreover, it is shown in the bivariate case in Theorem 6 of Bracco et al. [2019] that $I_{\Delta}({\bf z},X_{\bf z})$ behaves like $O(h^{-2})$ as ${\bf z}$ approaches a fault and $h=\hbox{diam}(X_{z})\rightarrow 0$, while it is bounded if ${\bf z}$ stays sufficiently far from all faults. In other words, the indicator tends to take large values at points close to faults and bounded values at points far from them, which justifies classifying as close to a fault the points of the set
${\cal F}(\alpha,X):=\{{\bf z}\in X:\,I_{\Delta}({\bf z},X_{\bf z})>\alpha\},$
@@
Thanks to these properties of the indicator, a reasonable choice of $\alpha$ is the median value of $I_{\Delta}$ computed on the nodes in $X$. This works well, except when $f$ is constant or linear in large subareas of the domain, since $\Delta f({\bf z})=0$ and so $I_{\Delta}({\bf z},X_{\bf z})\approx 0$ for any ${\bf z}$ belonging to these areas, and $\alpha$, chosen as the median value, would be very close to 00 and then all the points where the value of the indicator is slightly above 0 would be marked as close to a fault even if they are not. For this reason, we use the following two-step fault detection algorithm, that has shown excellent performance in extensive numerical tests on scattered data presented in Bracco et al. [2019].
The set of fault nodes ${\cal F}(X)$.
@@
    Set $\alpha_{1}=C_{1}\operatorname{median}\{I_{\Delta}({\bf z},X_{\bf z}):\,{\bf z}\in X\}$, and compute ${\cal F}(\alpha_{1},X)$.
@@
    Set $\alpha_{2}=C_{2}\operatorname{median}\{I_{\Delta}({\bf z},X_{z}):\,{\bf z}\in{\cal F}(\alpha_{1},X)\}$, and compute the final set of fault nodes as ${\cal F}(X):={\cal F}(\alpha_{2},{\cal F}(\alpha_{1},X))$.
Note that we use $C_{1}=1$ and $C_{2}=2$ in all numerical experiments in Section 6, the same values as in almost all experiments of Bracco et al. [2019].
@@
We now describe our approach to choosing the factor $\mu_{i}$ in (14) when adaptive fault-based viscosity is used. As input we need the set of nodes $X$, the values $U(t,x_{i})$ for all $x_{i}\in X$, and the spacing $h$ for the node set $X$. Spacing $h$ is a parameter that we use in the node generation as explained in Section 6, with $h\approx(|\Omega|/N)^{1/d}$. In addition we need three user specified parameters: $\mu$, the maximum factor for the artificial viscosity terms; the number $n_{\cal F}$ of nearest nodes for the computation of the fault indicator, which we choose to be the same number $n_{\cal F}=n_{\min}$ as in Section 3; and the constant $C_{3}$ that governs the transition zone between $\mu_{i}>0$ and $\mu_{i}=0$ in order to avoid abrupt changes in $\mu_{i}$. We use $C_{3}=5$ in the experiments of Section 6.
@@
    For each ${\bf z}={\bf x}_{i}\in X$, choose the set $X_{\bf z}=X_{i}^{init}$ consisting of $n_{\bf z}=n_{\cal F}$ nearest neighbors of ${\bf x}_{i}$ in $X$, and compute ${\cal F}(X)\subset X$ according to the above fault detection algorithm, with $f$ on $X$ given by $f({\bf x}_{j})=U(t,{\bf x}_{j})$, $j=1,\ldots,N$.
@@ 5 Summary of algorithms
As mentioned above, adding small artificial viscosity is motivated by the principle of vanishing viscosity, and its goal is it to allow accurate computation of the solution in the presence of shocks. As a side effect, solutions computed with artificial viscosity are smoother, which is normally not desired, but sometimes also produces better shapes by reducing non-physical behavior of the numerical solution near discontinuities. This latter positive effect is not significant for our method because consistent positive schemes satisfy the local maximum principle (9) and therefore do not suffer from artificial oscillations.
@@
Finally, we consider a scheme where viscosity is only added for points that are close to discontinuities (faults) detected by the method described in Section 4. The goal is to improve approximation quality in the smooth regions because the resulting discretization of $D_{{\boldsymbol{\eta}}_{i}}$ in (5) will allow a smaller set of influence $X_{i}$ and smaller seminorm $|{\bf w}_{i}|_{2,2}$ of the weight vector ${\bf w}_{i}=[w_{ij}]_{{\bf x}_{j}\in X_{i}}$ minimized in (25), in comparison to the weights $\tilde{w}_{ij}=w_{ij}-\mu_{i}v_{ij}$ effectively used in (14) when $\mu_{i}>0$. In addition, this scheme avoids the computation of the positive weights for the approximation of the Laplacian in most of the domain, leading to a significant computational saving. Note that the computation of the approximation of the Laplacian involved in the expression of the fault indicator (28) is inexpensive since a minimal differentiation formula without sign constraints on the weights is employed on a small fixed set of $n_{\cal F}$ neighbors, without an adaptive expansion of the influence set needed for the inequality constrained weights.
@@ 6 Numerical examples
The parameters of our schemes are chosen in a unified way in all examples. We always take $n_{\min}=10$, $n_{\max}=100$, $\Delta t=0.2h$ and choose $T$ such that $K=T/\Delta t$ is an integer. The viscosity or maximum viscosity factor is always $\mu=0.5h$, and additional parameters of the fault-based viscosity are chosen as $n_{\cal F}=10,C_{1}=1,C_{2}=2,C_{3}=5$.
@@
with respect to a reference solution $U_{\rm ref}$ given either by an exact solution (available in Example 1), or by a high resolution solution, which we obtain on a dense Cartesian grid by an appropriate higher order method provided in the code supplement of the book Hesthaven [2018]. In the case of an irregular set $X$, the values $U_{\rm ref}(T,{\bf x}_{i})$ are obtained by piecewise linear interpolation of $U_{\rm ref}(T,\cdot)$ on the Delaunay triangulation of the Cartesian grid. Note that $E_{1}$ and $E_{2}$ may be considered as discrete replacements for the $L^{1}$ and $L^{2}$ errors on $\Omega$, respectively.
@@ 6.1 Example 1: Inviscid Burgers’ equation with discontinuous initial condition
We choose spacing $h=0.01$ and generate $10000=1/h^{2}$ Halton points in the unit square, of which we remove 99 with distance to the boundary at most 0.0025. In addition, on the boundary 280 gridded nodes with step size $1/70$, and 4 nodes for each corner at distances $h/2$ and $h$ from it are added, see Figure 6.1. The total number of nodes in $X$ is therefore $N=10197$, very close to the number 10201 of nodes in a Cartesian grid with step size $h=0.01$.
Figure 6.1: Example 1: Irregular discretization nodes (only the south-west quarter of the domain is shown): interior nodes obtained by the Halton quasi-random sequence are in blue, boundary nodes in red color.
At each time step we identify inflow boundary nodes as those ${\bf x}_{i}\in X\cap\partial\Omega$, for which the approximated flow direction ${\boldsymbol{\eta}}_{i}={\bf F}^{\prime}(U(t,{\bf x}_{i}))$ points inside $\Omega$. For these nodes, we set $U(t+\Delta t,{\bf x}_{i}):=u(t+\Delta t,{\bf x}_{i})$ according to the above formula for the exact solution. For all other boundary and interior nodes the value $U(t+\Delta t,{\bf x}_{i})$ is computed according to the respective algorithm of Section 5 in use. Since the minimization problem (24) turns out to be infeasible even with $n_{\max}=100$ neighbors for many boundary nodes, we always set $\mu_{i}=0$ if ${\bf x}_{i}\in\partial\Omega$ in order to switch off the artificial viscosity for these nodes. For certain nodes (fewer than 20) situated in the immediate neighborhood of the inflow boundary, the minimization problem (25) is infeasible, and we use minimal numerical differentiation without inequality constraints to compute the weights $w_{ij}$. This does not generate instability in our experiments.
Numerical solutions at time $T=0.5$ obtained by Algorithms 13 are presented in Figures 6.2 (3D view) and 6.3 (view from above), in comparison to the exact solution. We observe that there are no instabilities in either solution, which is justified by the maximum principle (9) guaranteed for any positive scheme. However, the positive scheme without viscosity (Algorithm 1) completely misses an essential feature of the solution by straightening out one of discontinuity curves, which confirms that the stability and consistency of the scheme on their own do not guarantee a physically correct approximation when shocks are present. The shapes are much closer to the exact solution for Algorithms 2 and 3, indicating that the artificial viscosity gives a crucial contribution to the accuracy of the method. Moreover, the comparison between the solutions obtained by the positive scheme with constant and adaptive viscosity, see (c) and (d) in both Figures 6.2 and 6.3, reveals that the adaptive method significantly improves the solution in the crucial areas near discontinuities that look much sharper in the plots. In addition we provide errors $E_{1}$ and $E_{2}$ for the three solutions in Table 6.1. While they are significantly higher for Algorithm 1, there is little difference between Algorithms 2 and 3. However, this quantitative comparison is somewhat misleading as it is slightly in favor of Algorithm 2, in contrast to the clear visual evidence giving a significant advantage to Algorithm 3.
@@
(b) No viscosity (Algorithm 1).
(c) Constant viscosity (Algorithm 2).
(d) Adaptive viscosity (Algorithm 3).
Figure 6.2: Example 1: Visual comparison of numerical solutions for $T=0.5$.
@@
method $E_{1}$-error $E_{2}$-error no viscosity (Algorithm 1) 1.18e-01 3.89e-01 constant viscosity (Algorithm 2) 6.13e-02 2.30e-01 adaptive viscosity (Algorithm 3) 6.80e-02 2.73e-01

### v2

@@ 4 Adaptive fault-based viscosity
For each ${\bf x}_{i}\in X$, let $X_{i,\cal F}\subseteq X$ be the set of $n_{\cal F}$ points ${\bf x}_{j}\in X$ closest to ${\bf x}_{i}$. We define the fault indicator
$I_{\Delta}({\bf x}_{i},n_{\cal F}):=\frac{|\hat{\Delta}f({\bf x}_{i})|}{\displaystyle\sum_{{\bf x}_{j}\in X_{i,\cal F}}|w_{j}|\,\|{\bf x}_{j}-{\bf x}_{i}\|^{2}_{2}}\,,$ (29)
@@
$\hat{\Delta}f({\bf z}):=\sum_{{\bf x}_{j}\in X_{i,\cal F}}w_{j}f({\bf x}_{j})$ (30)
is the $\ell_{2}$-minimal differentiation formula approximating the Laplacian $\Delta f({\bf x}_{i})$ with polynomial exactness order $q=3$ and minimizing the seminorm $|{\bf w}|_{2,3}$. Note that the second fault indicator introduced in Bracco et al. [2019], employing the gradient instead of the Laplace operator, can also be used and gives comparable results in numerical experiments. We opted for the Laplacian-based one because it is cheaper. Indeed, the approximation of the scalar-valued Laplace operator by a numerical differentiation formula requires significantly less computational effort than approximating the vector-valued gradient operator.
@@
$I_{\Delta}({\bf x}_{i},n_{\cal F})\leq|f|_{1,1,\Omega_{loc}},$ (31)
with $\Omega_{loc}$ as in (22). Although (31) is stated in Bracco et al. [2019] only for the bivariate case, it is easy to extend it to any number of variables $d$. Indeed, this follows from the fact that Theorem 3 in Bracco et al. [2019] holds for any $d$ with the same proof. It is remarkable that the estimate (31) does not contain any unknown constants and therefore allows for comparison between different ${\bf x}_{i}$ independently of how irregular each local node set $X_{i,\cal F}$ is. Therefore, a large value of $I_{\Delta}({\bf x}_{i},n_{\cal F})$ indicates that $f$ is likely to be discontinuous in $\Omega_{loc}$, or at least its gradient tends to change abruptly there, that is the behavior expected from the solution of a conservation law in the vicinity of a shock front. Moreover, it is shown in the bivariate case in Theorem 6 of Bracco et al. [2019] that $I_{\Delta}({\bf x}_{i},n_{\cal F})$ behaves like $O(h^{-2})$ as ${\bf x}_{i}$ approaches a fault and $\hbox{diam}(X_{i,\cal F})\rightarrow 0$, while it is bounded if ${\bf x}_{i}$ stays sufficiently far from all faults. In other words, the indicator tends to take large values at points close to faults and bounded values at points far from them, which justifies classifying as close to a fault the points of the set
${\cal F}(\alpha,X):=\{{\bf x}_{i}\in X:\,I_{\Delta}({\bf x}_{i},n_{\cal F})>\alpha\},$
@@
Thanks to these properties of the indicator, a reasonable choice of $\alpha$ is the median value of $I_{\Delta}$ computed on the nodes in $X$. This works well, except when $f$ is constant or linear in large subareas of the domain, since $\Delta f({\bf z})=0$ and so $I_{\Delta}({\bf x}_{i},n_{\cal F})\approx 0$ for any ${\bf x}_{i}$ belonging to these areas, and $\alpha$, chosen as the median value, would be very close to 00. As a result, all the points where the value of the indicator is slightly above 0 would be marked as close to a fault even if they are not. For this reason, we use the following two-step fault detection algorithm, that has shown excellent performance in extensive numerical tests on scattered data presented in Bracco et al. [2019].
Computation of the set of fault nodes ${\cal F}(X)$.
@@
    Set $\alpha_{1}=C_{1}\operatorname{median}\{I_{\Delta}({\bf x}_{i},n_{\cal F}):\,{\bf x}_{i}\in X\}$, and compute ${\cal F}(\alpha_{1},X)$.
@@
    Set $\alpha_{2}=C_{2}\operatorname{median}\{I_{\Delta}({\bf x}_{i},n_{\cal F}):\,{\bf x}_{i}\in{\cal F}(\alpha_{1},X)\}$, and compute the final set of fault nodes as ${\cal F}(X):={\cal F}(\alpha_{2},{\cal F}(\alpha_{1},X))$.
Note that in all numerical experiments of Section 6 we use $C_{1}=1$ and $C_{2}=2$, the same values that proved to be effective in the experiments of Bracco et al. [2019], as our goal is to provide a robust choice, avoiding a fine-tuning of the parameters example by example.
@@
We now describe our approach to choosing the factor $\mu_{i}$ in (14) when adaptive fault-based viscosity is used. As input we need the set of nodes $X$, the values $U(t,{\bf x}_{i})$ for all ${\bf x}_{i}\in X$, and the spacing $h$ for the node set $X$. Spacing $h$ is a parameter that we use in the node generation as explained in Section 6, with $h\approx(|\Omega|/N)^{1/d}$. In addition we need three user specified parameters: $\mu$, the maximum factor for the artificial viscosity terms; the number $n_{\cal F}$ of nearest nodes for the computation of the fault indicator, which we choose to be the same number $n_{\cal F}=n_{\min}$ as in Section 3; and the constant $C_{3}$ that governs the transition zone between $\mu_{i}>0$ and $\mu_{i}=0$ in order to avoid abrupt changes in $\mu_{i}$, that otherwise have an adverse effect on the performance of the method. Note that the advantages of a continuous transition of the amount of the artificial viscosity between zero in the regions where the solution is smooth and its maximum value where the solution is discontinuous have been observed for other numerical methods, see for example Section 12.2.3 in Hesthaven [2018]. We use $C_{3}=5$ in the experiments of Section 6.
@@
    For each ${\bf x}_{i}\in X$, choose the set $X_{i,\cal F}=X_{i}^{init}$ consisting of $n_{\cal F}$ nearest neighbors of ${\bf x}_{i}$ in $X$, and compute ${\cal F}(X)\subset X$ according to the above fault detection algorithm, with $f$ on $X$ given by $f({\bf x}_{j})=U(t,{\bf x}_{j})$, $j=1,\ldots,N$.
@@
Fault points and the values of $\mu_{i}$ are illustrated in Figure 6.8 in the numerical examples section of the paper.
@@ 5 Summary of algorithms
As mentioned above, adding small artificial viscosity is motivated by the principle of vanishing viscosity, and its goal is to allow accurate computation of the solution in the presence of shocks. As a side effect, solutions computed with artificial viscosity are smoother, which is normally not desired, but sometimes also produces better shapes by reducing non-physical behavior of the numerical solution near discontinuities. This latter positive effect is not significant for our method because consistent positive schemes satisfy the local maximum principle (9) and therefore do not suffer from artificial oscillations.
@@
Finally, we consider a scheme where viscosity is only added for points that are close to discontinuities (faults) detected by the method described in Section 4. The goal is to improve approximation quality in the smooth regions because the resulting discretization of $D_{{\boldsymbol{\eta}}_{i}}$ in (5) will allow a smaller set of influence $X_{i}$ and smaller seminorm $|{\bf w}_{i}|_{2,2}$ of the weight vector ${\bf w}_{i}=[w_{ij}]_{{\bf x}_{j}\in X_{i}}$ minimized in (26), in comparison to the weights $\tilde{w}_{ij}=w_{ij}-\mu_{i}v_{ij}$ effectively used in (14) when $\mu_{i}>0$. In addition, this scheme avoids the computation of the positive weights for the approximation of the Laplacian in most of the domain, leading to a significant computational saving. In fact, the computation of the approximation of the Laplacian involved in the expression of the fault indicator (29) is less expensive since a minimal differentiation formula without sign constraints on the weights is employed, without an adaptive expansion of the influence set needed for the inequality constrained weights.
@@ 6 Numerical examples
The parameters of our schemes are chosen in a unified way in all examples. In particular, we make the viscosity factor $\mu$ directly and the time step $\Delta t$ inversely proportional to the maximum size of the characteristic velocity ${\bf F}^{\prime}(u_{0})$ by setting $\mu=0.5h\,v_{0}$ and $\Delta t=0.2h\,/v_{0}$, where $v_{0}:=\max_{{\bf x}\in\bar{\Omega}}\|{\bf F}^{\prime}(u_{0})\|_{\infty}$. We always take $n_{\min}=10$, $n_{\max}=100$, $\Delta t=0.2h$ and choose the final time $T$ such that $K=T/\Delta t$ is an integer. The viscosity or maximum viscosity factor is always $\mu=0.5h$, and additional parameters of the fault-based viscosity are chosen as $n_{\cal F}=10,C_{1}=1,C_{2}=2,C_{3}=5$.
@@
with respect to a reference solution $U_{\rm ref}$ given either by an exact solution (available in Example 1), or by a high resolution solution, obtained on a dense Cartesian grid with an appropriate higher order method provided in the code supplement of the book Hesthaven [2018]. In the case of an irregular set $X$, the values $U_{\rm ref}(T,{\bf x}_{i})$ are obtained by piecewise linear interpolation of $U_{\rm ref}(T,\cdot)$ on the Delaunay triangulation of the Cartesian grid. Note that $E_{1}$ and $E_{2}$ may be considered as discrete replacements for the $L^{1}$ and $L^{2}$ errors on $\Omega$, respectively.
@@ 6.1 Example 1: Inviscid Burgers’ equation with discontinuous initial condition
[…]
At each time step we identify inflow boundary nodes as those ${\bf x}_{i}\in X\cap\partial\Omega$, for which the approximated flow direction ${\boldsymbol{\eta}}_{i}={\bf F}^{\prime}(U(t,{\bf x}_{i}))$ points inside $\Omega$. For these nodes, we set $U(t+\Delta t,{\bf x}_{i}):=u(t+\Delta t,{\bf x}_{i})$ according to the above formula for the exact solution. For all other boundary and interior nodes the value $U(t+\Delta t,{\bf x}_{i})$ is computed according to the respective algorithm of Section 5 in use. Since the minimization problem (25) turns out to be infeasible even with $n_{\max}=100$ neighbors for many boundary nodes, we always set $\mu_{i}=0$ if ${\bf x}_{i}\in\partial\Omega$ in order to switch off the artificial viscosity for these nodes. For certain nodes (fewer than 20) situated in the immediate neighborhood of the inflow boundary, the minimization problem (26) is infeasible, and we use minimal numerical differentiation without inequality constraints to compute the weights $w_{ij}$. This does not generate instability in our experiments. In Figure 6.1(b) we show three examples of employed influence sets.
@@
(b) Halton nodes and influence sets.
Figure 6.1: Example 1: Exact solution and a zoom of the node set, with three highlighted influence sets $X_{i}$ (the encircled point is ${\bf x}_{i}$) of size $10$ (in cyan), $12$ (in magenta), and $15$ (in red).
Numerical solutions at time $T=0.5$ obtained by Algorithms 13 are presented in Figures 6.2 and 6.3. Here, Figure 6.2 gives 3D visualizations of the discrete solutions on the left and point cloud-based scatter plots over a transparent exact solution on the right, and Figure 6.3 provides the view from above, with the plot in 6.3(a) based on the exact solution on the same Halton nodes as in Figure 6.2. Note that the wrinkles near discontinuities in the left plots of Figures 6.2 are visualization artefacts, whereas the actual solutions are discrete as shown on the right. We observe that there are no instabilities in either solution, which is justified by the maximum principle (9) guaranteed for any positive scheme. However, the positive scheme without viscosity (Algorithm 1) completely misses an essential feature of the solution by straightening out one of discontinuity curves, which confirms that the stability and consistency of the scheme on their own do not guarantee a physically correct approximation when shocks are present. The shapes are much closer to the exact solution for Algorithms 2 and 3, indicating that the artificial viscosity gives a crucial contribution to the accuracy of the method. Moreover, the comparison between the solutions obtained by the positive scheme with constant and adaptive viscosity, see (c) and (d) in both Figures 6.2 and 6.3, reveals that the adaptive method significantly improves the solution in the crucial areas near discontinuities that look much sharper in the plots. In addition we provide errors $E_{1}$ and $E_{2}$ for the three solutions in Table 6.1. While they are significantly higher for Algorithm 1, there is little difference between Algorithms 2 and 3. However, this quantitative comparison is somewhat misleading as it is slightly in favor of Algorithm 2, in contrast to the clear visual evidence giving a significant advantage to Algorithm 3.
(a) No viscosity (Algorithm 1).
(b) Constant viscosity (Algorithm 2).
(c) Adaptive viscosity (Algorithm 3).
Figure 6.2: Example 1: Visual comparison of numerical solutions for $T=0.5$. For each algorithm, the 3D visualization of the discrete solution (left) is shown together with the point cloud-based scatter plot over a transparent exact solution (right).
@@
method $E_{1}$-error $E_{2}$-error no viscosity (Algorithm 1) 1.18e-01 3.89e-01 constant viscosity (Algorithm 2) 6.43e-02 2.38e-01 adaptive viscosity (Algorithm 3) 7.04e-02 2.78e-01